the area of a 6cm by 4m rectangle is to be doubled. by what equal amount must each dimension be increased
Answer by mananth(16949)

Area of rectangle = 6*4 = 24 m^2
.
(let it be increased by x m
(x+6)(x+4)=48
x^2+10x+24=48
X^2+10x-24=0
X^2+12x-2x-24=0
x(x+12)-2(x+12)=0
(x+12)(x-2)=0
x= 2 m
Ignore the negative value of x=-12